OpenStudy (anonymous):

What is the mass in mg of a 2.275mL sample, if it has a density of 3.75kg/L ?

OpenStudy (abmon98):

Density=mass/volume to convert 1 kilogram 1 000 000 milligrams A milligram is a unit of mass in the Metric System. The symbol for milligram is mg. There are 1,000 milligrams in a gram. What is a kilogram (kg)? A kilogram is a unit of mass in the Metric System. The symbol for kilogram is kg. There are 1,000 grams in a kilogram. Change 2.275 ml to L 1 ml =0.001 L 2.275=x L x=0.00275 L x/0.00275=3.75*10^6

OpenStudy (abmon98):

x in the last equation represent the mass

OpenStudy (anonymous):

I don't think that's correct. The answer from what you supplied is 10312.5 and that's not correct according to the website it was entered on.

OpenStudy (anonymous):

mass=volume*density you just regulate these units.

OpenStudy (joannablackwelder):

@Abmon98 was correct but made a small typo. The volume is 0.002275 mL, not 0.00275.

OpenStudy (anonymous):

(2.275mL) x (3.75 x 10^3 mg/mL) = 8531 mg